Here is the more information about the market for onions in Vancouver .



Price per unit- 5,8,11,14,17etc



Quantity supplied-7,16,25,34,43etc



*Find the equation for the supply curve .



*What is the quality supplied when price =16?



* What is the cost of producing quantity supplied = 10 ?

1
Expert's answer
2022-01-18T09:58:59-0500

Coordinates=(7,5), (16,8), (25,11), (34,14), (43,17)

Use any two coordinates to find the Slope

Form of a straight line= Y=X+mX

Where; m= Slope

X= Quantity supplied

Y= Price

"\frac{(8-5)}{(16-7)}=\frac {3}{9}=\frac{1}{3}"

Using the Slope

Assume a coordinate with (X,Y) and use it with any of the above coordinates

"\frac{(Y-8)}{(X-16)}=\frac{1}{3}"

"(Y-8)3=X-16"

3Y-24=X-16

3Y=X+8

"Y=\frac{1}{3}X+\frac{8}{3}"


"\therefore Q_s=\frac{1}{3}p+\frac{8}{3}"

Quantity supplied at price 16

"Q_s=\frac{1}{3}(16)+\frac{8}{3}"

"=\frac{16}{3}+\frac{8}{3}=\frac{24}{3}= 8"


Cost of producing quantity supplied= "8\times 10= 80"